OpenStudy (anonymous):

3x^2 + 22x - 16 = 0 solve by factoring

OpenStudy (anonymous):

I am relatively good at these but I am stuck.

OpenStudy (campbell_st):

the easy way is multiply 3 and - 16 gives -48.... find the factors of -48 that add to 22 -2 and 24 then its \[\frac{(3x - 2)(3x + 24)}{3} = \frac{3(3x -2)(x + 8)}{3}\] cancel the common factor of 3 leaves (3x -2)(x + 8)

OpenStudy (anonymous):

3x^2+24x-2x-16=0 3x(x+8)-2(x+8)=0 (3x-2)(x+2)=0

OpenStudy (anonymous):

thank you both so much. I am really grateful

OpenStudy (anonymous):

(3x-2)(x+*)

OpenStudy (anonymous):

x+8

OpenStudy (anonymous):

Wait, I need to solve to zero, I keep coming up with 2/3 and -8 its not right I don't think. Can you help me finish it off?

OpenStudy (anonymous):

what answer is given for this

OpenStudy (campbell_st):

the solution of 2/3 and -8 is correct

OpenStudy (anonymous):

when I check it, it comes out all wacky . Maybe it's just me

OpenStudy (anonymous):

could we go through that part together please? If you have time

OpenStudy (anonymous):

yua that ans is correct

OpenStudy (campbell_st):

ok... 3(-8)^2 + 22(-8) - 16 = 192 - 176 - 16 = 0 and 3(2/3)^2 + 22(2/3) - 16 = 4/3 + 44/3 - 16 = 48/3 - 16 = 0 both solutions work

OpenStudy (anonymous):

W.O.W. I feel silly, I had simply forgotten to square. Thank you so much for your help.

OpenStudy (campbell_st):

hope it helps